The specific heat capacity of ethanol is 2440J/kg °C. How many joules of energy will be required to heat 150g ethanol to 35°C if the starting temperature is 22°C?


1
Expert's answer
2021-08-26T08:10:35-0400

We have to consider that all the thermal energy Q that has to be required can be found if we consider that the mass of ethanol m has a specific heat capacity C. By using the equation that relates these variables with the final and initial temperatures (Tf and Ti) we can find the energy "Q=m_{ethanol}C_{ethanol}\Delta T". After substitution we find:


"Q=m_{ethanol}C_{ethanol} (T_f-T_i)
\\ Q=(150 \,g)(\frac{1\,kg}{1000\,g})(2440 \frac{J}{kg\, ^\circ C}) (35-22)^\circ C
\\ Q=4758\,J"


In conclusion, the energy required to heat 150 g of ethanol from 22 to 35 °C is 4758 Joules.


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.